# Mirror Online Conformal Prediction with Intermittent Feedback

### v2

@@ Abstract
Online conformal prediction enables the runtime calibration of a pre-trained artificial intelligence model using feedback on its performance. Calibration is achieved through set predictions that are updated via online rules so as to ensure long-term coverage guarantees. While recent research has demonstrated the benefits of incorporating prior knowledge into the calibration process, this has come at the cost of replacing coverage guarantees with less tangible regret guarantees based on the quantile loss. This work introduces intermittent mirror online conformal prediction (IM-OCP), a novel runtime calibration framework that integrates prior knowledge, while maintaining long-term coverage and achieving sub-linear regret. IM-OCP features closed-form updates with minimal memory complexity, and is designed to operate under potentially intermittent feedback.
@@ I Introduction
IN safety-critical and high-stakes sequential decision-making processes, such as in robotics [1, 2], wireless communications [3, 4], finance [5], and medicine [6, 7], it is crucial to accurately quantify uncertainty in order to reliably predict the potential outcomes of given actions. Producing reliable uncertainty estimates in these scenarios is often challenging due to the non-stationarity of the data-generating distribution and to the possibility that feedback on the quality of the predictions may be only intermittently available [8].
A widely used approach to quantify uncertainty complements point predictions with prediction sets that satisfy coverage guarantees [9]. Notably, adaptive conformal inference (ACI) [10] converts black-box predictions into prediction sets that enjoy distribution-free, long-term coverage guarantees. This is done by adaptively adjusting prediction set sizes based on feedback about past decisions [11, 12, 13, 14, 15]. Intermittent feedback can be accommodated via intermittent ACI (I-ACI) [16].
In contrast to ACI, the recently introduced Bayesian ACI (B-ACI) adopts a data-centric Bayesian objective that incorporates not only feedback, but also prior information on the prediction scores [17]. When the prior information is well specified, B-ACI can yield prediction sets with lower regret in terms of the quantile loss as compared to ACI. A lower regret with respect to the quantile loss implies better performance in stochastic settings with independent data [17]. However, B-ACI does not guarantee the key calibration requirement of deterministic long-term coverage. Furthermore, it requires storing past data, as well as solving a convex problem at each calibration step.
As summarized in Table I, existing online calibration schemes either cannot incorporate prior information or sacrifice long-term coverage guarantees. To address these limitations, we introduce intermittent mirror online conformal prediction (IM-OCP). IM-OCP builds on online mirror descent (OMD) [18, 19, 20] to integrate prior knowledge, while employing an importance weighting strategy to handle intermittent feedback [21]. This design enables IM-OCP to retain the computational and memory efficiency of ACI, while incorporating prior information. We showcase the merits of IM-OCP for the task of received signal strength indicator (RSSI)-based localization under intermittent feedback [22, 23, 24].
@@ II-A Online Calibration via Set Prediction
We consider the problem of online calibration in sequential decision-making problems. At each round $t$, a learner observes an input $X_{t}\in\mathcal{X}$ and aims to predict the corresponding label $Y_{t}\in\mathcal{Y}$. To this end, the predictor assigns a negatively oriented score $s(X_{t},Y)$ to each value $Y\in\mathcal{Y}$, which is assumed to be upper bounded by a constant $B<\infty$, i.e., $s(X,Y)\leq B$. For example, the quadratic loss is defined as $s(X_{t},Y)=|f(X_{t})-Y|^{2}$, where $f(X_{t})$ is the prediction. Using predictor’s output, a prediction set $\mathcal{C}_{t}$ is produced that includes all labels $Y\in\mathcal{Y}$ with a score no larger than a threshold $r_{t}$, i.e.,
@@
Given a target miscoverage rate $\alpha\in[0,1]$, the goal of calibration is to leverage the outlined feedback information to optimize, in an online fashion, the thresholds $\{r_{t}\}_{t=1}^{T}$ in (1) so as to ensure the long-term coverage guarantee over the time horizon $T$ [10]
@@
While the condition (2) applies to an arbitrary sequence $\{(X_{t},Y_{t})\}_{t=1}^{T}$, in some settings one may be justified to assume that the sequence consists of independent and identically distributed (i.i.d.) samples. In this case, the probabilistic coverage condition $\Pr\{Y_{t}\notin\mathcal{C}_{t}\}=\alpha$ can be met for each time $t$ by selecting the threshold $r_{t}$ as the $(1-\alpha)$-quantile of the score distribution. This quantile can be estimated by minimizing the empirical quantile loss
$\ell_{1-\alpha}(r,r^{*}_{t})=(\alpha-\mathbbm{1}\{r<r^{*}_{t}\})(r-r^{*}_{t})$ (3)
as $q_{\alpha}(r^{*}_{1:T})=\min_{u\in\mathbb{R}}\sum_{t=1}^{T}\ell_{1-\alpha}(u,r^{*}_{t})$. In fact, with i.i.d. data, the estimate $q_{\alpha}(r^{*}_{1:T})$ tends to the true $(1-\alpha)$-quantile in the limit $T\rightarrow\infty$.
To assess the capacity of a calibration procedure to perform well also in the case of i.i.d. data, it is thus useful to evaluate the extent to which the thresholds $\{r_{t}\}_{t=1}^{T}$ deviate from the estimated quantile $q_{\alpha}(r^{*}_{1:T})$. Note that the latter can be only evaluated in hindsight, i.e., at time $T$, while the thresholds $r_{t}$ must be produced online at time $t$. This deviation is measured by the regret
@@ II-B Adaptive Conformal Inference
$r_{t}$ $=r_{t-1}-\eta_{t-1}\nabla\ell_{1-\alpha}(r_{t-1},r_{t-1}^{*})$
$=r_{t-1}-\eta_{t-1}(\alpha-E_{t-1}),$ (5)
where $\eta_{t}>0$ is the step size. Thanks to the regret guarantees of OGD and the restorative nature of the quantile loss [25], given a learning rate sequence $\eta_{t}=c/\sqrt{t}$ with $c>0$ independent of $T$, ACI enjoys both sublinear regret and long-term coverage guarantee (2) with $\gamma=1/2$[14].
@@ II-C Bayesian Adaptive Conformal Inference
ACI and I-ACI do not allow the integration of prior knowledge about the data distribution in the calibration process. In contrast, B-ACI assumes the availability of a prior distribution $P(r)$, with support $[0,B]$, on the scores $\{r^{*}_{t}\}_{t\geq 1}$. Leveraging this information, along with score feedback, at each time step $t$, the threshold $r_{t}$ is obtained by solving the convex optimization problem
$r_{t}=\arg\min_{r}\left\{\eta_{t}\psi(r)+\frac{1-\eta_{t}}{t-1}\sum_{i=1}^{t-1}\ell_{1-\alpha}(r,r_{i}^{*})\right\},$ (7)
where the convex regularizing function $\psi(r)=\mathbb{E}_{r^{*}\sim P}[\ell_{1-\alpha}(r,r^{*})]$ accounts for prior knowledge.
Unlike ACI and I-ACI, which have constant memory complexity and a simple update rule that can be computed in closed form, B-ACI requires linear memory to store the sequence $\{r^{*}_{t}\}_{t\geq 1}$ of feedback signals, and it requires to solve a convex minimization problem at each round $t$. B-ACI attains sublinear regret, potentially outperforming ACI when the prior knowledge is well specified. However, B-ACI is not guaranteed to satisfy the long-term coverage condition (2).
@@
$r_{t}=\arg\min_{r}\left\{\eta_{t}\psi(r)+\frac{1-\eta_{t}}{t-1}\sum_{i=1}^{t-1}\ell_{1-\alpha}(r,r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}\right\}.$ (8)
In Appendix -A, we show that IB-ACI achieves sub-linear expected regret (see (14)).
III Intermittent Mirror Online Conformal Prediction
Given the state of the art summarized in Table I, this paper introduces IM-OCP, a novel prior-dependent calibration scheme with intermittent feedback that exhibits low complexity, sublinear regret, and coverage guarantees. IM-OCP tackles the problem of online calibration via an OMD strategy [18, 19, 20]. This way, IM-OCP retains the simplicity and performance guarantees of first-order methods such as ACI, while allowing for the incorporation of prior information.
For a given prior distribution $P(r)$ defined in $[0,B]$ for the scores $r^{*}$ and a parameter $\sigma>0$, IM-OCP introduce a regularizing function $R(\cdot)$ of the form
$R(r)=\mathbb{E}_{r^{*}\sim P}[\ell_{1-\alpha}(r,r^{*})]+\frac{\sigma}{2}|r|^{2},$ (9)
where the first term follows B-ACI, while the second term ensures the strong convexity of the function $R(\cdot)$. In Appendix -B, we show that the regularizing function (9) is closed, differentiable, $\mu$-strongly convex for some real number $\mu>\sigma$, and $L$-smooth for some real number $L\leq\max_{r\in[0,B]}\{P(r)\}+\sigma$.
Given the regularizing function $R(r)$, the associated Bregman divergence is denoted as $\mathcal{B}_{R}(u,v)=R(u)-R(v)-(\nabla R(v))^{T}(u-v)$. The corresponding mirror map is given by $\mathcal{M}(\cdot)=\nabla R(\cdot)$, with its inverse map defined as $\mathcal{M}^{-1}(\cdot)=(\nabla R)^{-1}(\cdot)$.
At each time $t$, IM-OCP updates the threshold via the OMD rule
$r_{t}=\mathcal{M}^{-1}\left(\mathcal{M}(r_{t-1})-\eta_{t-1}(\alpha-E_{t-1})\frac{\text{obs}_{t-1}}{p_{t-1}}\right).$ (10)
When the feedback $E_{t-1}$ is not observed, and thus we have $\text{obs}_{t-1}=0$, the update (10) corresponds to the identity map. In contract, when the feedback $E_{t-1}$ is observed, the update (10) amounts to a gradient descent step in the dual space associated with the mirror map.
III-A Theoretical Guarantees
We now analyze the performance of IM-OCP, showing that it satisfies expected long-term coverage while also enjoying sublinear expected regret. Proofs are deferred to Supplementary Material.
@@ III-A Theoretical Guarantees
Similarly to [14], we first establish the following ancillary result stating that the iterates produced by IM-OCP are bounded.
@@
We now prove that IM-OCP enjoys sublinear regret. Similar to (11), we define the IM-OCP regret as
@@
where $D_{T}=\max_{t\in\{1,\dots,T\}}\mathcal{B}_{R}(q_{\alpha}(r^{*}_{1:T}),r_{t})$.
@@ IV-A Setting
We consider a multi-building, multi-floor indoor localization problem and adopt the UJIIndoorLoc dataset [26]. The goal is to localize a device positioned in one of three buildings based on the signal power received from $M=520$ access points [22, 23, 24]. Accordingly, the input $X_{t}\in\mathbb{R}^{M}$ corresponds to an $M$-dimensional vector of RSSI values, while the label $Y_{t}=[Y_{t}^{\mathrm{Log}},Y_{t}^{\mathrm{Lat}}]^{T}\in\mathbb{R}^{2}$ represents the longitude and latitude coordinates of the corresponding device to be localized. We apply online calibration to a pre-trained localization model $\hat{Y}_{i}=f_{\text{ELM}}(X_{i})$, implemented using an extreme learning machine with a hidden layer of size 256 [24]. The model is trained using a training set consisting of 1000, 2000, and 8000 samples randomly selected from Buildings 1, 2, and 3, respectively. Since the accuracy of $f_{\text{ELM}}(\cdot)$ is expected to increase in the building index $i$, due to the increasingly larger datasets as $i$ grows, we set a decreasing probability of receiving feedback as $p_{1}=0.5$ for the first building, $p_{2}=0.3$ for the second, and $p_{3}=0.1$ for the third.
@@ IV-B Results
We perform online calibration of the base predictor using I-ACI, B-ACI, and IM-OCP, considering both triangular and truncated Gaussian prior distributions $P(r)$ defined in the interval $[0,B]$. Specifically, for the triangular prior distribution, the mode is set to 0.1, while for the truncated Gaussian prior distribution, the mean is set to 0.1 and the variance to 2. For all schemes, we consider a held-out data sequence of length $T=2400$, with samples randomly selected from each building with equal probability, and set the target miscoverage rate to $\alpha=0.1$.
@@
In Fig. 1, we present the miscoverage rate as a function of time index $t$. Both IM-OCP and ACI, which satisfy the miscoverage guarantee (2), exhibit a miscoverage rate that converges to the target value $\alpha$. In contrast, IB-ACI, which does not come with a theoretical coverage guarantee, obtains a final miscoverage rate higher than the target value.
@@
In Fig. 2, we report the cumulative quantile loss for all algorithms. By leveraging prior information, IM-OCP achieves a lower cumulative loss—and consequently, a smaller regret—compared to I-ACI, while remaining competitive with B-ACI.
Overall, these results demonstrate that IM-OCP satisfies the coverage guarantees while incorporating prior information and achieving a lower regret compared to I-ACI.

### v3

@@ Abstract
Online conformal prediction enables the runtime calibration of a pre-trained artificial intelligence model using feedback on its performance. Calibration is achieved through set predictions that are updated via online rules so as to ensure long-term coverage guarantees. While recent research has demonstrated the benefits of incorporating prior knowledge into the calibration process, this has come at the cost of replacing coverage guarantees with less tangible regret guarantees based on the quantile loss. This work introduces intermittent mirror online conformal prediction (IM-OCP), a novel runtime calibration framework that integrates prior knowledge, while maintaining long-term coverage and achieving sublinear regret. IM-OCP features closed-form updates with minimal memory complexity, and is designed to operate under potentially intermittent feedback.
@@ I Introduction
IN safety-critical and high-stakes sequential decision-making processes, such as in robotics [1, 2], wireless communications [3, 4], finance [5], and medicine [6, 7], it is important to accurately quantify uncertainty in order to reliably predict the potential outcomes of given actions. Producing reliable uncertainty estimates in these scenarios is often challenging due to the non-stationarity of the data-generating distribution and to the possibility that feedback on the quality of the predictions may be only intermittently available [8].
A formal approach to quantify uncertainty complements point predictions with prediction sets that satisfy coverage guarantees [9]. Notably, online conformal prediction (OCP) [10] schemes convert predictions produced by any black-box model into prediction sets that enjoy distribution-free, long-term coverage guarantees. Specifically, adaptive conformal inference (ACI) [11] adjusts prediction set sizes based on feedback about past decisions [11, 12, 13, 14, 15, 16]. ACI can be wrapped around any predictive model, including classifiers [17], regressors [11], and sequential models (large language models) [18]. Sparse feedback can be accommodated via intermittent ACI (I-ACI) [19].
In contrast to ACI, the recently introduced Bayesian ACI (B-ACI) is an OCP method that adopts a data-centric Bayesian objective incorporating not only feedback, but also prior information on the prediction scores [20]. When the prior information is well specified, B-ACI can yield prediction sets with lower regret in terms of the quantile loss as compared to ACI, which implies better performance in stochastic settings with independent data [20]. However, B-ACI does not guarantee deterministic long-term coverage like ACI. Furthermore, it requires storing past data, as well as solving a convex problem at each calibration step.
As summarized in Table I, existing OCP schemes either cannot incorporate prior information or sacrifice long-term coverage guarantees (see Sec. II for details). To address these limitations, this work introduces intermittent mirror OCP (IM-OCP), a novel online calibration framework, that integrates prior knowledge, while ensuring long-term deterministic coverage guarantee and exhibiting constant memory and low-complexity updates like ACI. IM-OCP builds on online mirror descent (OMD) [21, 22, 23], inheriting from OMD also sublinear regret with respect to the quantile loss. We showcase the merits of IM-OCP for the task of received signal strength indicator (RSSI)-based localization under intermittent feedback [24, 25, 26].
@@ II-A Online Calibration via Set Prediction
We consider the problem of online calibration in sequential decision-making problems. At each round $t$, a learner observes an input $X_{t}\in\mathcal{X}$ and aims to predict the corresponding label $Y_{t}\in\mathcal{Y}$. To this end, the predictor assigns a negatively oriented score $s(X_{t},Y)$ to each value $Y\in\mathcal{Y}$, which is assumed to be upper bounded by a constant $B<\infty$, i.e., $s(X,Y)\leq B$. For example, the quadratic loss is defined as $s(X_{t},Y)=|f(X_{t})-Y|^{2}$, where $f(X_{t})$ is the prediction. Using predictor’s output, a prediction set $\mathcal{C}_{t}$ is produced that includes all labels $Y\in\mathcal{Y}$ with a score no larger than a threshold $r_{t}$, i.e., [10]
@@
Given a target miscoverage rate $\alpha\in[0,1]$, the goal of calibration is to leverage the outlined feedback information to optimize, in an online fashion, the thresholds $\{r_{t}\}_{t=1}^{T}$ in (1) so as to ensure the long-term deterministic coverage guarantee over the time horizon $T$ [11]
@@
While the condition (2) applies to an arbitrary sequence $\{(X_{t},Y_{t})\}_{t=1}^{T}$, in some settings one may be justified to assume that the sequence consists of independent and identically distributed (i.i.d.) samples. In this case, the probabilistic coverage condition $\Pr\{Y_{t}\notin\mathcal{C}_{t}\}=\alpha$ can be met for each time $t$ by selecting the threshold $r_{t}$ as the $(1-\alpha)$-quantile of the score distribution. This quantile can be estimated by minimizing the empirical average of the quantile loss [12, 13, 14, 15, 16]
$\ell_{1-\alpha}(r,r^{*}_{t})=(\alpha-\mathbbm{1}\{r<r^{*}_{t}\})(r-r^{*}_{t}),$ (3)
as we have equality $q_{\alpha}(r^{*}_{1:T})=\min_{u\in\mathbb{R}}\sum_{t=1}^{T}\ell_{1-\alpha}(u,r^{*}_{t})$. In fact, with i.i.d. data, the estimate $q_{\alpha}(r^{*}_{1:T})$ tends to the true $(1-\alpha)$-quantile in the limit $T\rightarrow\infty$.
To assess the capacity of a calibration procedure to perform well also in the case of i.i.d. data, it is thus useful to evaluate the extent to which the thresholds $\{r_{t}\}_{t=1}^{T}$ deviate from the estimated quantile $q_{\alpha}(r^{*}_{1:T})$. Note that the latter can be only evaluated in hindsight, i.e., at time $T$, while the thresholds $r_{t}$ must be produced online at time $t$. This deviation is measured by the regret [21, 22, 23]
@@ II-B Adaptive Conformal Inference
$r_{t}=r_{t-1}-\eta_{t-1}(\alpha-E_{t-1}),$ (5)
where $\eta_{t}>0$ is the step size. By setting $\eta_{t}=c/\sqrt{t}$, ACI achieves a sublinear regret $\mathcal{O}(B\sqrt{T})$, while also ensuring the coverage guarantee in (2) with $\gamma=1/2$ [15]. Furthermore, given the single-pass nature of OGD [21, 22, 23], ACI enjoys constant memory complexity $\mathcal{O}(1)$.
@@ II-C Bayesian Adaptive Conformal Inference
ACI and I-ACI do not allow the integration of prior knowledge about the data distribution in the calibration process. In contrast, B-ACI [20] assumes the availability of a prior distribution $P(r)$, with support $[0,B]$, on the scores $\{r^{*}_{t}\}_{t\geq 1}$. Leveraging this information, along with score feedback, at each time step $t$, the threshold $r_{t}$ is obtained by solving the convex optimization problem
$r_{t}=\arg\min_{r}\left\{h_{t}\psi(r)+\sum_{i=1}^{t-1}\ell_{1-\alpha}(r,r_{i}^{*})\right\},$ (7)
where $h_{t}=\frac{\eta_{t}(t-1)}{1-\eta_{t}}$ and the convex regularizing function $\psi(r)=\mathbb{E}_{r^{*}\sim P}[\ell_{1-\alpha}(r,r^{*})]$ accounts for prior knowledge.
B-ACI attains a sublinear regret $\mathcal{O}(D\sqrt{T})$, where the constant $D$ can be smaller than $B$ when the prior distribution is well specified [20]. Thus, B-ACI has the potential to outperform ACI. However, B-ACI is not guaranteed to satisfy the long-term coverage condition (2). Furthermore, B-ACI requires a memory growing linearly over time as $\mathcal{O}(T)$ to store the sequence of feedback signals $\{r^{*}_{t}\}_{t\geq 1}$, and it necessitates solving the convex problem in (7) at each round $t$.
@@
$r_{t}=\arg\min_{r}\left\{h_{t}\psi(r)+\sum_{i=1}^{t-1}\ell_{1-\alpha}(r,r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}\right\}.$ (8)
In Appendix -A, we show that IB-ACI achieves sublinear expected regret (see (14)).
III Intermittent Mirror Online Conformal Prediction (IM-OCP)
In this section, we present IM-OCP, a novel prior-dependent calibration scheme with intermittent feedback that exhibits low complexity, sublinear regret, and coverage guarantees.
III-A Algorithm Development
IM-OCP incorporates prior information via a regularization term, and updates threshold $r_{t}$ via the application of OMD.
III-A1 Regularizer Encoding Prior Information
For a given prior distribution $P(r)$ on the scores $r^{*}$ defined on the interval $[0,B]$, IM-OCP introduces a regularizer $R(\cdot)$ of the form
$R(r)=\psi(r)+\frac{\sigma}{2}|r|^{2},$ (9)
where the first term follows B-ACI, while the second term ensures the strong convexity of the function $R(\cdot)$ for any fixed $\sigma>0$. In Appendix -B, we show that the function $R(\cdot)$ is closed, differentiable, $\mu$-strongly convex for some real number $\mu>\sigma$, and $L$-smooth for some real number $L\leq\max_{r\in[0,B]}\{P(r)\}+\mu$.
III-A2 Threshold Update via OMD
Using the regularizer $R(r)$ in (9), IM-OCP updates the threshold at each time $t$ according to the OMD rule [21, 22, 23]. To describe it, define the mirror map induced by function $R(r)$ as $\mathcal{M}_{R}(\cdot)=\nabla R(\cdot)$, and the corresponding inverse as $\mathcal{M}_{R}^{-1}(\cdot)=(\nabla R)^{-1}(\cdot)$. This mirror map $\mathcal{M}_{R}$ captures the geometry induced by the prior information. The update rule of IM-OCP is then given as
$r_{t}=\mathcal{M}_{R}^{-1}\left(\mathcal{M}_{R}(r_{t-1})-\eta_{t-1}(\alpha-E_{t-1})\frac{\text{obs}_{t-1}}{p_{t-1}}\right).$ (10)
When the feedback $E_{t-1}$ is observed ($\text{obs}_{t-1}$$=$$1$), this update corresponds to a gradient step in the dual space defined by function $R(\cdot)$. This adjusts the threshold $r_{t}$ in a direction informed by both the observed feedback and the prior knowledge embedded in the mirror map $\mathcal{M}_{R}$. In contrast, when no feedback is observed ($\text{obs}_{t-1}$$=$00), the update in (10) reduces to the identity map, and the threshold remains unchanged.
III-B Theoretical Guarantees
We now present the theoretical guarantees of IM-OCP. While the regret performance follows with minor changes from the standard theory of OMD [21, 22, 23], the analysis of long-term coverage is novel and challenging, requiring new proofs around the boundedness of the iterates produced by the update rule (10). Proofs are deferred to the supplementary material.
@@ III-B Theoretical Guarantees
We first establish the following ancillary result stating that the iterates produced by IM-OCP are bounded.
@@
IM-OCP also enjoys sublinear regret. Similar to (11), we define the IM-OCP regret as
@@
where $D_{T}=\max_{t\in\{1,\dots,T\}}\mathcal{B}_{R}(q_{\alpha}(r^{*}_{1:T}),r_{t})$, with the $\mathcal{B}_{R}(u,v)=R(u)-R(v)-(\nabla R(v))^{T}(u-v)$.
@@ IV-A Setting
We consider a multi-building, multi-floor indoor localization problem and adopt the UJIIndoorLoc dataset [27]. The goal is to localize a device positioned in one of three buildings based on the signal power received from $M=520$ access points [24, 25, 26]. Accordingly, the input $X_{t}\in\mathbb{R}^{M}$ corresponds to an $M$-dimensional vector of RSSI values, while the label $Y_{t}=[Y_{t}^{\mathrm{Log}},Y_{t}^{\mathrm{Lat}}]^{T}\in\mathbb{R}^{2}$ represents the longitude and latitude coordinates of the corresponding device to be localized. We apply online calibration to a pre-trained localization model $\hat{Y}_{i}=f_{\text{ELM}}(X_{i})$, implemented using an extreme learning machine with a hidden layer of size 256 [26]. The model is trained using a training set consisting of 1000, 2000, and 8000 samples randomly selected from Buildings 1, 2, and 3, respectively. Since the accuracy of $f_{\text{ELM}}(\cdot)$ is expected to increase in the building index $i$, due to the increasingly larger datasets as $i$ grows, we set a decreasing probability of receiving feedback as $p_{1}=0.5$ for the first building, $p_{2}=0.3$ for the second, and $p_{3}=0.1$ for the third. In these results, we focus on longitude prediction, using the prediction set $\mathcal{C}_{t}^{\text{Lon}}(X_{t},r_{t}^{\text{Lon}}):=\left[[f_{\text{ELM}}(X_{i})]_{1}-r_{t}^{\text{Lon}},[f_{\text{ELM}}(X_{i})]_{1}+r_{t}^{\text{Lon}}\right]$. A similar approach can be applied to latitude prediction.
@@ IV-B Results
We perform online calibration of the base predictor using I-ACI, IB-ACI, and IM-OCP, considering both triangular and truncated Gaussian prior distributions $P(r)$ defined in the interval $[0,B]$. Specifically, for the triangular prior distribution, the mode is set to 0.1, while for the truncated Gaussian prior distribution, the mean is set to 0.1 and the variance to 2. For all schemes, we consider a held-out data sequence of length $T=2400$, with samples randomly selected from each building with equal probability, and set the target miscoverage rate to $\alpha=0.1$.
@@
In Fig. 1, we present the miscoverage rate as a function of the time index $t$. Confirming the expected miscoverage guarantees (13) provided in Theorem 1, the proposed IM-OCP exhibits a miscoverage rate that converges to the target level $\alpha$. In contrast, IB-ACI, which does not come with a theoretical coverage guarantee, obtains a final miscoverage rate higher than the target value.
In Fig. 2, we report the cumulative quantile loss for all algorithms. By leveraging prior information, IM-OCP achieves a lower cumulative loss—and consequently, a smaller regret—compared to I-ACI, while remaining competitive with B-ACI. Overall, these results verify the theoretical coverage guarantees of IM-OCP, while also demonstrating its capacity to incorporate prior information and achieve a lower regret compared to I-ACI.